OpenStudy (dariyana):

PLEASE HELP S = 4v^2 for v. will the answer be s/4=v^2

OpenStudy (dariyana):

@cwrw238

OpenStudy (anonymous):

\[S=4v^2 \to \frac{ S }{ 4 } = v²\]Using square root on both sides,\[\sqrt{\frac{ S }{ 4 }} = \sqrt{v²} \to \frac{ \sqrt{S} }{ 2 } = v\]

OpenStudy (dariyana):

i am confused why would i square root

OpenStudy (dariyana):

@zairhenrique

OpenStudy (anonymous):

I did it to get the value of v. I think v² = s/4 still correct.

OpenStudy (dariyana):

ok so you just did it to get rid of the squared on the v

OpenStudy (anonymous):

Yeah!

OpenStudy (dariyana):

ok thank you so much

OpenStudy (anonymous):

you welcome :]